# Primes and polygonal numbers

### v3

@@ Abstract.
For all integers $r,s>2$, each linear combination of an $r$-gonal number and an $s$-gonal number with coprime positive integer coefficients produces infinitely many primes. However, for any given pair of coprime positive integer coefficients of the polygonal numbers, the sum of the reciprocals of such primes diverges if and only if $r=s=4$.
@@ 1. Introduction and the main result
A prime is a sum of two squares if and only if it is of the form $14k+$ 1. Since there are infinitely many primes of the form $14k+$ 1, it follows that there are infinitely many primes which are sum of two squares. Since the sums of couples of triangular numbers are intimately related with the sums of couples of squares, viz. $n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares, it is natural to expect that the primes which are sums of two triangular numbers will also have an equally appealing description in terms of some arithmetic progressions. However, it turns out not to be the case. In fact, the sum of the reciprocals of such primes converges. One may still wonder whether there are infinitely many such primes or more generally, whether for a given pair of integers $r,s>2$, there are infinitely many primes which are sum of an $r$-gonal number and an $s$-gonal number, where the $m$-th $r$-gonal number is defined by (see [2])
$T(r,m):=(r-2)\frac{m(m-1)}{2}+m.$ (1.1)
Even more generally, one may ask whether for each pair of integers $r,s>2$ and for every pair of mutually coprime positive integers $a$ and $b$, there are infinitely many primes of the form $aT(r,m)+bT(s,n).$ We answer this in the affirmative.
@@
From the above theorem, via Abel’s summation formula, we obtain
@@
Given $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, the sum of the reciprocals of the primes of the form $aT(r,m)+bT(s,n)$ diverges if and only if $r=s=4$.
In particular, Theorem 1 solves an open problem in [1], which asks how many primes are sums of two triangular numbers.
@@
In fact, there are infinitely many such primes in each of the residue classes ±$\pm 1$ modulo $4$.
@@
Given two mutually coprime positive integers $a$ and $b$, the number of the primes in each of the residue classes ±$\pm 1$ modulo $4$, which are less than $N$ and which are linear combinations of two odd-sided polygonal numbers with coefficients $a$ and $b$ is
@@
Since a number $n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares (see [1]) and since a prime is a sum of two squares if and only if it is of the form $14k+$ 1 (see [8]), we have
Corollary 3.
The number of primes $p\leq N$ such that both $p$ and $14p+$ 1 are a sums of two squares is
@@
By an admissible arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms. Now, looking at Theorem 2, one may wonder whether every admissible arithmetic progression contains infinitely many primes which are sums of two triangular numbers. However, that does not hold in general! For example, if $n$ is congruent to $5$ or $8$ modulo 9, then $14n+$ 1 is not a sum of two squares. So, none of the arithmetic progressions $\{9k+5\}_{k\in\mathbb{N}}$ and $\{9k+8\}_{k\in\mathbb{N}}$ contains a sum of two triangular numbers. Further, Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply the following.
@@
$\{aT(r,m)+bT(s,n)\}_{m,n\in\mathbb{N}}$
misses infinitely many primes from every admissible arithmetic progression.
In other words, one can only describe the primes of the form $am^{2}+bn^{2}$ by a finite set of congruences for certain $a,b\in\mathbb{N}$ (see [3]). However, for any pair of $a,b\in\mathbb{N}$, if we replace either $m^{2}$ or $n^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers fails to produce infinitely many primes contained in any given admissible arithmetic progression.
@@
Length Arithmetic progressions
@@
10 73091, 135461, 197831, 260201, 322571, 384941, 447311, 509681, 572051, 634421
The above table suggests the existence of arbitrarily long arithmetic progressions among the primes which are sums of two triangular number. More generally, we ask whether there are arbitrarily long arithmetic progressions among the primes which are linear combinations of two polygonal numbers with coprime positive integer coefficients.
Open Question 1.
For each choice of $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, are there are arbitrarily long arithmetic progressions among the primes of the form $aT(r,m)+bT(s,n)?$
@@ 2. Proofs of the theorems
The main ingredients of the proof of Theorem 1 are a couple of results by Iwaniec [5], which we state below.
@@
A Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be of an irreducible polynomial in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$. Then
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\gg\frac{N}{\log N},$ (2.1)
@@
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\asymp\frac{N}{(\log N)^{3/2}},$ (2.2)
@@
B Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g\in\mathbb{Z}[x,y]$. Then $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly dependent if and only if
@@
Proof of Theorem 1: For a pair of integers $r,s\geq 3$ and for a pair of mutually coprime integers $\mu,\nu>0$, considering without loss of generality all possible parities of the natural numbers $m$ and $n$ in
$\mu T(r,m)+\nu T(s,n),$
@@
Since $\mu,\nu>0$ and $r,s>2$, the first two equalities imply that $aa^{\prime}>0$ and $bb^{\prime}>0$ and the last equality implies that there exists $k\in\mathbb{Q}$ such that $\frac{a}{b}=-\frac{a^{\prime}}{b^{\prime}}=k$. Hence, substituting $kb$ for $a$ and $-kb^{\prime}$ for $a^{\prime}$ in the first equality and dividing both sides of it with the corresponding sides of the second equality above, we obtain
@@
$\pi_{i}(N):=\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{i}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1.$
@@ Acknowledgments
All the computations for Table 1 was done using PARI/GP. Also, Bhattacharya thankfully acknowledges the support from the Department of Science and Technology, Govt. of India through the INSPIRE Faculty Award IFA17-MA104 and Rahaman thankfully acknowledges the support from the Ministry of Education, Govt. of India through the Prime Minister Research Fellowship 0501972.

### v4

@@ Abstract.
A linear combination $aT_{r}(m)+bT_{s}(n)$ of an $r$-gonal number $T_{r}(m)$ and an $s$-gonal number $T_{s}(n)$ with mutually coprime positive integer coefficients $a$ and $b$ produces infinitely many primes as $m$ and $n$ varies over the natural numbers, whereas the sum of the reciprocals of such primes converges unless $T_{r}(m)=m^{2}$ and $T_{s}(n)=n^{2}$. For each pair of coprime positive integers $a$ and $b$, there are arbitrary long arithmetic progressions among the primes of the form $am^{2}+bn^{2}$.
@@ 1. Introduction and the main result
A prime is a sum of two squares if and only if it is of the form $4k+1$. Since there are infinitely many primes of the form $4k+1$, it follows that there are infinitely many primes which are sum of two squares. Since the sums of couples of triangular numbers are intimately related with the sums of couples of squares, viz. $n$ is a sum of two triangular numbers if and only if $4n+1$ is a sum of two squares, it is natural to expect that the primes which are sums of two triangular numbers will also have an equally appealing description in terms of some arithmetic progressions. However, it turns out not to be the case. In fact, the sum of the reciprocals of such primes converges. One may still wonder whether there are infinitely many such primes or more generally, whether for a given pair of integers $r,s>2$, there are infinitely many primes which are sum of an $r$-gonal number and an $s$-gonal number, where the $m$-th $r$-gonal number is defined by (see [2])
$T_{r}(m)\;:=\;(r-2)\frac{m(m-1)}{2}+m.$ (1.1)
Even more generally, one may ask whether for each pair of integers $r,s>2$ and for every pair of mutually coprime positive integers $a$ and $b$, there are infinitely many primes of the form $aT_{r}(m)+bT_{s}(n).$ We answer this in the affirmative.
@@
Recall that $f\asymp g$ if and only if there exists a pair of constants $C,C^{\prime}>0$ such that $Cg\leq f\leq C^{\prime}g$. In the first case above, we have $T_{4}(m)=m^{2}$. The surprise is in the other case, where even if one of the summands is a multiple of a square, the corresponding set of primes has a lower rate of growth. From the above theorem, via Abel’s summation formula, we obtain
@@
Given $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, the sum of the reciprocals of the primes of the form $aT_{r}(m)+bT_{s}(n)$ diverges if and only if $r=s=4$.
In particular, Theorem 1 answers a question from [1], which asks how many primes are sums of two triangular numbers.
@@
In fact, there are infinitely many such primes in each of the residue classes $\pm 1$ modulo $4$.
@@
Given two mutually coprime positive integers $a$ and $b$, the number of the primes in each of the residue classes $\pm 1$ modulo $4$, which are less than $N$ and which are linear combinations of two odd-sided polygonal numbers with coefficients $a$ and $b$ is
@@
In particular, for $a=b=1$ and for the primes congruent to $1$ modulo $4$, we have the following.
Corollary.
For each odd integer $r>1$, the number of primes $p\leq N$, which are sums of two $r$-gonal numbers and also sums of two squares, is
$\asymp\frac{N}{(\log N)^{3/2}}.$
Since a number $n$ is a sum of two triangular numbers if and only if $4n+1$ is a sum of two squares*** It follows from (1.1) that for all $r$, we have $4(r-2)(T_{r}(m)+T_{r}(n))+(r-4)^{2}\;=\;\left((r-2)(m-n)\right)^{2}+\left((r-2)(m+n)-r+4\right)^{2}.$ Conversely, if $4(r-2)M\;+\;(r-4)^{2}\;=\;A^{2}\;+\;B^{2}$ such that $A\;\pm\;B\;\equiv\;r\pmod{2(r-2)}$, then $M$ is a sum of two $r$-gonal numbers. (see [1]) and since a prime is a sum of two squares if and only if it is of the form $4k+1$ (see [8]), it follows that
Corollary.
The number of primes $p\leq N$ such that both $p$ and $4p+1$ are sums of two squares is
@@
By a primitive arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms. Now, looking at Theorem 2, one may wonder whether every primitive arithmetic progression contains infinitely many primes which are sums of two triangular numbers. However, that does not hold in general! For example, if $n$ is congruent to $5$ or $8$ modulo 9, then $4n+1$ is not a sum of two squares. So, none of the arithmetic progressions $\{9k+5\}_{k\in\mathbb{N}}$ and $\{9k+8\}_{k\in\mathbb{N}}$ contains a sum of two triangular numbers. This prompts us to refine our query as follows:
Question 1.
Given $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, if there is no congruence obstruction, does every primitive arithmetic progression contain infinitely many primes which are of the form $aT_{r}(m)+bT_{s}(n)?$
Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply the following.
@@
$\{aT_{r}(m)+bT_{s}(n)\}_{m,n\in\mathbb{N}}$
misses infinitely many primes from every primitive arithmetic progression.
In other words, one can only describe the primes of the form $am^{2}+bn^{2}$ by a finite set of congruences for certain $a,b\in\mathbb{N}$ (see [3]). However, for any pair of $a,b\in\mathbb{N}$, if we replace either $m^{2}$ or $n^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers fails to produce infinitely many primes contained in any given primitive arithmetic progression.
@@
Length
Arithmetic progressions
@@
10
73091, 135461, 197831, 260201, 322571, 384941, 447311,
509681, 572051, 634421
The following table suggests the existence of arbitrarily long arithmetic progressions among the primes which are sums of two triangular number. More generally, we ask whether there are arbitrarily long arithmetic progressions among the primes which are linear combinations of two polygonal numbers with coprime positive integer coefficients.
Question 2.
For each choice of $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, are there are arbitrarily long arithmetic progressions among the primes of the form $aT_{r}(m)+bT_{s}(n)?$
@@ 2. Proofs of the theorems
The main ingredients of the proof of Theorem 1 are a couple of results by Iwaniec [5], which we state below. A
@@
Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be an irreducible polynomial in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$. Then
$\Bigg{(}\sum_{\begin{subarray}{c}\text{prime\hskip 3.0pt}p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\text{\hskip 3.0ptfor $x,y\in\mathbb{N}$}\end{subarray}}1\Bigg{)}\;\;\gg\;\;\frac{N}{\log N},$ (2.1)
@@
$\Bigg{(}\sum_{\begin{subarray}{c}\text{prime\hskip 3.0pt}p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P(x,y)\text{\hskip 3.0ptfor $x,y\in\mathbb{N}$}\end{subarray}}1\Bigg{)}\;\;\asymp\;\;\frac{N}{(\log N)^{3/2}},$ (2.2)
@@
B Let $P(x,y)\;=\;ax^{2}+bxy+cy^{2}+ex+fy+g$. Then $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly dependent if and only if
@@
Proof of Theorem 1. For a pair of integers $r,s\geq 3$ and for a pair of mutually coprime integers $\mu,\nu>0$, considering without loss of generality all possible parities of the natural numbers $m$ and $n$ in
$\mu T_{r}(m)+\nu T_{s}(n),$
@@
Since $\mu,\nu>0$ and $r,s>2$, the first two equalities imply that $aa^{\prime}>0$ and $bb^{\prime}>0$ and the last equality implies that there exists $k\in\mathbb{Q}$ such that $\frac{a}{b}=-\frac{a^{\prime}}{b^{\prime}}=k$. Hence, substituting $kb$ for $a$ and $kb^{\prime}$ for $a^{\prime}$ in the first equality and dividing both sides of it with the corresponding sides of the second equality above, we obtain
@@
$\pi_{i}(N)\;:=\;\sum_{{\begin{subarray}{c}\text{prime\hskip 3.0pt}p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{i}(x,y)\text{\hskip 3.0ptfor $x,y\in\mathbb{N}$}\end{subarray}}}1.$
@@ Acknowledgments
The authors would like to thank Prof. Don Zagier for his comments and suggestions. All the computations for Table 1 was done using PARI/GP. Also, Bhattacharya thankfully acknowledges the support from the Department of Science and Technology, Govt. of India through the INSPIRE Faculty Award IFA17-MA104 and Rahaman thankfully acknowledges the support from the Ministry of Education, Govt. of India through the Prime Minister Research Fellowship 0501972.